How do you find m and b for the following problem: 3x-2y=5
Answer by nyc_function(2645)


The y-intercept form for an equation is y = mx + b, where m is the slope and b is the y-intercept.
We need to change your equation into the form y = mx + b.
To do so, we isolate y on the left side of the equation.
3x - 2y = 5
-2y = -3x + 5
y = (-3x + 5)/-2

y = (3/2)(x) - 5/2
From this final equation, we know that m = 3/2 and b = -5/2